1. The position vector r of a missile at any time t is given by
r = ti + (8t − t
2
)j,
where i and j are unit vectors in the horizontal and upward vertical directions. Find its
(a) average velocity from t = 2 to t = 3,
(b) velocity, speed and direction of motion when t = 4,
(c) position vector when it is moving parallel to the vector i − 2j,
(d) acceleration vector.
2. The position vector r of a moving particle at time t after the start of the motion is given by
r = 5(1 + 4t)i + 5(19 + 2t − t
2
)j.
Find the initital velocity of the particle. At time t = T, the particle is moving at right angles to its initial
direction of motion. Find the value of T and the distance of the particle from its initial position at this
time.
1
Expert's answer
2020-04-30T19:14:05-0400

(a) average velocity from t = 2 to t = 3,


"r = ti + (8t \u2212 t^2)j"

"average\phantom{i}velocity=\frac{r(3)-r(2)}{3-2}\\"

"r(3) = 3i + (8(3) − (3)^2)j\\r(3) = 3i + 15j\\r(2)= 2i + (8(2) −(2)^2)j\\r(2) = 2i + 12j"


hence


"average\phantom{i}velocity=\frac{r(3)-r(2)}{3-2}\\average\phantom{i}velocity=\frac{(3i+15j)-(2i+12j)}{3-2}\\average\phantom{i}velocity=\frac{i+3j}{1}\\average\phantom{i}velocity=i+3j"


(b) velocity, speed and direction of motion when t = 4,

velocity function

"r = ti + (8t \u2212 t^2)j"

"velocity=\frac{dr}{dt}\\velocity=\frac{dr}{dt}=i+(8-2t)j"


hence at t=4

"velocity=i+0j\implies velocity=i"


for speed

"speed=|\frac{dr}{dt}|\\\frac{dr}{dt}=i"

hence speed

"speed=\sqrt{1^2}\implies speed=1"


direction of motion

"direction=\frac{v(t)}{|v(t)|}\\direction=\frac{i}{1}\\direction=i"

"tan\theta=\frac{y-component}{x-component}\phantom{i}of\phantom{i}the\phantom{i}velocity\\
\theta=tan^{-1}(\frac{0}{1})\\\theta=0^{0}"

both methods implies the particle is moving along the x-axis

c.) position vector when it is moving parallel to the vector i − 2j

"\begin{pmatrix}
 t \\
 8t-t^2
\end{pmatrix}=k\begin{pmatrix}
 1 \\
 -2
\end{pmatrix}\\\implies t=k\\8t-t^2=-2k\\"

solving the two equations simultaneously by substitution

"8t-t^2=-2t\\hence\\t=10
\\\therefore r=10i-20j"

d.) acceleration vector

"acceleration=\frac{dv}{dt}\\\frac{dv}{dt}=-2j"



2.) initial velocity

"r = 5(1 + 4t)i + 5(19 + 2t − t^2)j. \\velocity=\frac{dr}{dt}\\\frac{dr}{dt}=20i+5(2-2t)j\\velocity=20i+5(2-2t)j
\\at\phantom{i}t=0"

initial velocity

"v=20i-10tj"


value of T

"r = 5(1 + 4t)i + 5(19 + 2t \u2212 t^2)j"

two vectors are said to be perpendicular if their dot product is equal to zero

initial displacement is when t=0

"t=0\\r = 5i + 95j"

at t=T

"r = 5(1 + 4T)i + 5(19 + 2T \u2212 T^2)j"

then

"(5(1 + 4T)i + 5(19 + 2t − T^2)j).(5i+95j)=0\\(5(1 + 4T)5)+(5(19 + 2T− T^2)95)=0"

open the bracket and solve for T

"25+100T+9025+950T-475T^2=0\\475T^2-1050T-9050=0"

solving quadratic

T=5.6080

"distance\phantom{i}t=5.6080\\""\\r = 5(1 + 4t)i + 5(19 + 2t − t^2)j\\r=r = 5(1 + 4(5.6080))i + 5(19 + 2(5.6080) − (5.6080)^2)j\\
r=117.16i-6.16832j\\distance=\sqrt{(117.16)^2+(-6.16832)^2}\\distance=\sqrt{13764.51377}\\distance=117.322"